OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 2 choices: x=18 x=8 x=1 x=16

OpenStudy (anonymous):

log2*(x)-log2*(4)=2 is this your question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

there is something wrong

OpenStudy (anonymous):

hints. log_2 (a) + log_2 (b) = log_2 (ab) log_2 (a) - log_2 (b) = log_2 (a/b) also log_2 (32) = 5 is another way of writing 2^(5) = 32

OpenStudy (anonymous):

hmmmmmmmm

OpenStudy (anonymous):

\[\log_{2} x-\log_{2} 4=2\] you asking this

OpenStudy (anonymous):

\[\log_{2} (x/4)=2\]

OpenStudy (anonymous):

x/4=4 x=16

OpenStudy (anonymous):

\[\log_{2} x -\log_{2} 4 =2\] I think this is the question. if this ,ans is 8

OpenStudy (anonymous):

sorry 16